# Towards Understanding Deep Learning Model in Image Recognition via Coverage Test

### v1

@@ 1. Introduction
Due to their complex network structure, DNNs may produce different judgments for inputs with minor differences, leading to erroneous results. Therefore, testing DNNs to improve system quality has become extremely important. Existing work uses traditional software testing methods to generate test cases to check for potential vulnerabilities (Ma et al., 2018). To evaluate the effectiveness of these generated test cases, various coverage metrics have been proposed, such as neuron coverage (Ma et al., 2018) and Modified Condition/Decision Coverage (MC/DC) (Sun et al., 2019a). Also, to focus on the behavior of neurons in different locations within the network structure, various coverage metrics have different characteristics (Xie et al., 2022). For example, as shown in Figure 1, statement coverage mainly checks the executable statements involved in the test cases; neuron coverage evaluates the proportion of neurons covered by the test cases.
Many works have focused on the priority of testing inputs (Shi et al., 2021) to guide security personnel to achieve detection effects in a shorter time, thereby saving resource consumption during testing. Therefore, current empirical research focuses on comparing test case priority indicators (Shi et al., 2021) with gradient-based attack methods (Goodfellow et al., 2014) for effectiveness. These methods aim to select more suitable evaluation metrics to generate test cases to evaluate model quality (Aghababaeyan et al., 2023). However, these studies on neuron coverage only focus on test cases and lack research that considers the DNN model itself as an influencing factor. Nowadays, many excellent DNN structures have been proposed, such as LeNet, AlexNet, and VGG (Kriegeskorte and Golan, 2019). DNNs with different structures not only have different model structures but also different model depths. As a model component in AI models, the variation of DNNs is an important factor to be considered (Nguyen et al., 2020).
@@ 2.1. Neuron Network Coverage
In order to make structured coverage metrics more suitable for deep neural networks, a variety of neural network coverage metrics have been proposed to guide deep neural networks. The operation of software depends on the execution continuity. For example, traditional software depends on the previous statement, while deep neural networks depend on the previous layer of neurons. As shown in Figure 1, the traditional program has branches. After the program runs, the test case input can determine whether there is a problem in the code corresponding to the program statement. Therefore, the more code statements the traditional program covers, the higher the probability of finding code problems. For deep neural networks, the coverage should be based on the neuron activity to determine whether the input has passed through all possible cases.
According to the strength of neuronal activity, neural network coverage can be divided into main functional area coverage and boundary area coverage. Moreover, due to the different granularity features in different layers of deep neural networks (Li et al., 2021c), the layer coverage index is also introduced in this paper.
@@ 2.2. Main Functional Area Coverage Indicators
Neuron Coverage. Neuron Coverage (NC) (Pei et al., 2017) refers to the ratio of the number of neurons whose activation values exceed a threshold when the neural network is input with a test input set to the total number of all neurons. The formula for NC can be expressed as where represents a neuron, and all neurons constitute the set ; represents the test input, and all test inputs constitute the set ; represents the activation value of a neuron, means the activation result after inputting on neuron ; is the threshold for a neuron to be activated.
K-multisection Neuron Coverage. K-multisection Neuron Coverage (KMNC) (Ma et al., 2018) refers to dividing the activation value domain of all neurons into k equal parts, and the ratio of the part covered by the test input set to the total part when inputting the neural network. The activation value domain is obtained by inputting the training dataset into the model and observing the activation data statistics of all neurons during the training process. The formula for KMNC can be expressed as where represents the activation value of input instance in neuron ; is a value from 1 to ; represents the KMNC of neuron .
Neuron Boundary Coverage. The KMNC metric is applied to the activation value range , i.e., it can only cover in the activation value range. Therefore, the Neuron Boundary Coverage (NBC) metric (Ma et al., 2018) is defined as the ratio of the number of neurons covered outside the activation value range to the total number of neurons. The formula for the NBC metric can be represented as where is a neuron in (Shi et al., 2021), and the number of neurons that exist in the lower domain and the upper domain are both . In detail, we denote the size of each region as .
@@ 4. Future Work
Deep neural network models are widely used in feature extraction for many downstream tasks such as image segmentation, object detection, and pose estimation. With the wide application of deep neural networks, model security testing has also attracted a lot of attention (Li et al., 2024a; Kong et al., 2024; Niu et al., 2024; Liu and Li, 2025; Li et al., 2021a, 2024b; Zou et al., 2025; Mao et al., 2024; Bu et al., 2025b; Li et al., 2021b, 2017; Wang et al., 2024; Liu et al., 2024; Li et al., 2025, 2024c; Bu et al., 2025a). Therefore, in order to test the security of these models, various coverage metrics have been proposed to detect model units covered by input test cases (Sun et al., 2019a; Shi et al., 2021; Pei et al., 2017). The influencing factors of the coverage detection effect were explored to improve the evaluation accuracy of the test cases on the model. The main process of the coverage index in the model is improved to enhance the interpretability of the deep neural network. In summary, future research on DNN coverage may have the following directions.

### v2

@@ 1. Introduction
Due to their complex network structure, DNNs may produce different judgments for inputs with minor differences, leading to erroneous results (Ghaleb et al., 2023; Tsankov et al., 2018; Tikhomirov et al., 2018). Therefore, testing DNNs to improve system quality has become extremely important. Existing work uses traditional software testing methods to generate test cases to check for potential vulnerabilities (Ma et al., 2018). To evaluate the effectiveness of these generated test cases, various coverage metrics have been proposed, such as neuron coverage (Ma et al., 2018) and Modified Condition/Decision Coverage (MC/DC) (Sun et al., 2019a). Also, to focus on the behavior of neurons in different locations within the network structure, various coverage metrics have different characteristics (Xie et al., 2022). For example, as shown in Figure 1, statement coverage mainly checks the executable statements involved in the test cases; neuron coverage evaluates the proportion of neurons covered by the test cases.
Many works have focused on the priority of testing inputs (Shi et al., 2021) to guide security personnel to achieve detection effects in a shorter time, thereby saving resource consumption during testing (Bhargavan et al., 2016; Ferreira et al., 2020). Therefore, current empirical research focuses on comparing test case priority indicators (Shi et al., 2021) with gradient-based attack methods (Goodfellow et al., 2014) for effectiveness. These methods aim to select more suitable evaluation metrics to generate test cases to evaluate model quality (Aghababaeyan et al., 2023). However, these studies on neuron coverage only focus on test cases and lack research that considers the DNN model itself as an influencing factor (Zhong et al., 2024; White et al., 2016; Sajnani et al., 2016; Tian et al., 2022). Nowadays, many excellent DNN structures have been proposed, such as LeNet, AlexNet, and VGG (Kriegeskorte and Golan, 2019). DNNs with different structures not only have different model structures but also different model depths. As a model component in AI models, the variation of DNNs is an important factor to be considered (Nguyen et al., 2020).
@@ 2.1. Neuron Network Coverage
In order to make structured coverage metrics more suitable for deep neural networks, a variety of neural network coverage metrics have been proposed to guide deep neural networks (He et al., 2020; Ghaleb and Pattabiraman, 2020; Neamtiu et al., 2005; Yang et al., 2024). The operation of software depends on the execution continuity. For example, traditional software depends on the previous statement, while deep neural networks depend on the previous layer of neurons (Xiao et al., 2025; Wang et al., 2021; Qian et al., 2022; Chu et al., 2023). As shown in Figure 1, the traditional program has branches. After the program runs, the test case input can determine whether there is a problem in the code corresponding to the program statement (Nadini et al., 2021; Ante, 2023; Tolmach et al., 2021; Gao et al., 2020). Therefore, the more code statements the traditional program covers, the higher the probability of finding code problems. For deep neural networks, the coverage should be based on the neuron activity to determine whether the input has passed through all possible cases (Huang et al., 2021; Ma et al., 2024; Szabo, 1996; Ante, 2022).
According to the strength of neuronal activity, neural network coverage can be divided into main functional area coverage and boundary area coverage (Mikolov, 2013; Choromanski et al., 2020; Beltagy et al., 2020). Moreover, due to the different granularity features in different layers of deep neural networks (Li et al., 2021b), the layer coverage index is also introduced in this paper.
@@ 2.2. Main Functional Area Coverage Indicators
Neuron Coverage. Neuron Coverage (NC) (Pei et al., 2017) refers to the ratio of the number of neurons whose activation values exceed a threshold when the neural network is input with a test input set to the total number of all neurons (Srinivasa et al., 2020; Malkov and Yashunin, 2018; Ain et al., 2019). The formula for NC can be expressed as where represents a neuron, and all neurons constitute the set ; represents the test input, and all test inputs constitute the set ; represents the activation value of a neuron, means the activation result after inputting on neuron ; is the threshold for a neuron to be activated.
K-multisection Neuron Coverage. K-multisection Neuron Coverage (KMNC) (Ma et al., 2018) refers to dividing the activation value domain of all neurons into k equal parts, and the ratio of the part covered by the test input set to the total part when inputting the neural network (Liang et al., 2024; Zhu et al., 2022; Zhang, 2024; Pasqua et al., 2023). The activation value domain is obtained by inputting the training dataset into the model and observing the activation data statistics of all neurons during the training process. The formula for KMNC can be expressed as where represents the activation value of input instance in neuron ; is a value from 1 to ; represents the KMNC of neuron .
Neuron Boundary Coverage. The KMNC metric is applied to the activation value range , i.e., it can only cover in the activation value range (Bojanowski et al., 2017; Zhang et al., 2024). Therefore, the Neuron Boundary Coverage (NBC) metric (Ma et al., 2018) is defined as the ratio of the number of neurons covered outside the activation value range to the total number of neurons. The formula for the NBC metric can be represented as where is a neuron in (Shi et al., 2021), and the number of neurons that exist in the lower domain and the upper domain are both . In detail, we denote the size of each region as .
@@ 4. Future Work
Deep neural network models are widely used in feature extraction for many downstream tasks such as image segmentation, object detection, and pose estimation. With the wide application of deep neural networks, model security testing has also attracted a lot of attention (Li et al., 2024a; Kong et al., 2024; Niu et al., 2024; Liu and Li, 2025; Li et al., 2021a, 2024c; Zou et al., 2025; Mao et al., 2024; Bu et al., 2025b; Li and others, 2021; Li et al., 2017; Wang et al., 2024; Liu et al., 2024; Li et al., 2025, 2024b; Bu et al., 2025a). Therefore, in order to test the security of these models, various coverage metrics have been proposed to detect model units covered by input test cases (Sun et al., 2019a; Shi et al., 2021; Pei et al., 2017). The influencing factors of the coverage detection effect were explored to improve the evaluation accuracy of the test cases on the model. The main process of the coverage index in the model is improved to enhance the interpretability of the deep neural network. In summary, future research on DNN coverage may have the following directions.